OpenStudy (anonymous):

Choose the correct simplification of the expression (2b over c) ^3 2b^3 over c^3 8b^3 over c^3 2b^3 over c 2b over c^3

OpenStudy (nurali):

\[\frac{ 2^{3} b ^{1*3} }{ c ^{1*3}}\]

OpenStudy (anonymous):

doesnt seem to match any of the answer choices.

OpenStudy (anonymous):

\[\left(\frac{2b}{c}\right)^3\]?

OpenStudy (anonymous):

\[\frac{2^3b^3}{c^3}=\frac{8b^3}{c^3}\]

OpenStudy (nurali):

solve it

OpenStudy (anonymous):

2b over c^3 is what i think

OpenStudy (anonymous):

if i knew how to solve it, i wouldnt have posted.

OpenStudy (nurali):

when u follow or solve what i have given u will get the answer of second option.

OpenStudy (nurali):

if correct please click best response.